Find the ordered pair for the vertex.
y = 2x2 - 4x + 6

Answer by nyc_function(2741)

y = 2x^2 - 4x + 6
We need the point (x,y), which is our vertex.
To find x, use x = -b/2a, where b = -4 and a = 2 from your function.
x = -(-4)/2(2)
x = 4/4
x = 1
To find y, we plug x = 1 into your function and solve for y.
y = 2(1)^2 - 4(1) + 6
y = 2 - 4 + 6
y = -2 + 6
y = 4
The vertex is the point (1,4).
Done!